1/4(8y + 4) - 17 = -1/2(4y - 8) Would I clear out the fractions by four which is the lcm? 1(8y + 4) - 17 = - 2(4y - 8)
8 + 4 - 17 = -8y - 8
12-17= -16
- 5 = - 16??? I am so lost Can I have some help please?

Answer by rfer(16322)

2y+1-17=-2y+4
2y+2y=+4+16
4y=20
y=20/4
y=5
Multiply though by the fractions.
If it helps change them to decimals for the calculations.